OpenStudy (anonymous):

What are the intercepts of the equation? -3x+4y=-12

OpenStudy (jdoe0001):

to find the x-intercept, set y = 0, and solve for "x" to find the y-intercept, set x = 0, and solve for "y" keep in mind that an intercept is when the graph "intercepts" or "touches" the axis 

OpenStudy (anonymous):

Thank you very much. That helped much more then the first answer that i got lol. Could you tell me if I am correct on this? x-intercept: (-3, 0) y-intercept: (0, 4)

OpenStudy (jdoe0001):

\(\bf -3x+4y=-12 \implies -3(0)+4y=-12 \implies y = -3\quad y-intercept\\ (0,-3)\\ -3x+4y=-12\implies -3x+4(0)=-12 \implies x = 4\quad x-intercept\\ (4, 0)\)

OpenStudy (anonymous):

I see thank you :)